lilybean123:

A paper popcorn cone measures 4 inches in diameter and is 9 inches high. What is the volume of popcorn that can be contained within this cone?

jhonyy9:

so do you know formula for volume of a cone ?

lilybean123:

no lol

jhonyy9:

but without this not can calcule the volume of popcorn

lilybean123:

ok

Ruthy:

4/2=2 2*4*9=72cube

lilybean123:

that is not on the answer choices

jhonyy9:

@Ruthy do you can explain please what you ve wrote above ?

Ruthy:

Kk

Ruthy:

To get the radius u divide the diameter by 2 then 2is ur radius 2*4*9 the height= 72cube